Given the following price and dividend information, calculate the $1 invested equivalent. (Round to 4 decimals)

Year    Price    Dividend
0 50.72
1 43.54 1.75
2 49.22 2.10
3 51.30 2.20
4 52.45 2.50
5 56.35 2.75
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

C D E Dividend Return А в 1 2 Year 30 1 12 3 4 15 Price $50.72 $43.54 $49.22 $51.30 $52.45 $56.35 $1.75 $2.10 $2.20 $2.50 $2.

Cell reference -

Year Price Dividend Return 1.75 50.72 43.54 49.22 51.3 52.45 56.35 =((C4-C3)+D4)/C3 =((C5-C4)+D5)/C4 =((C6-C5)+D6)/C5 =((C7-C

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.